A company has a 12% WACC and is considering two mutually exclusive investments (that cannot be repeated) with the following cash flows:

0 1 2 3 4 5 6 7
Project A -$300 -$387 -$193 -$100 $600 $600 $850 -$180
Project B -$405 $132 $132 $132 $132 $132 $132 $0
  1. What is each project's NPV? Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to the nearest cent.

    Project A: $  

    Project B: $  

  2. What is each project's IRR? Do not round intermediate calculations. Round your answers to two decimal places.

    Project A:   %

    Project B:   %

  3. What is each project's MIRR? (Hint: Consider Period 7 as the end of Project B's life.) Do not round intermediate calculations. Round your answers to two decimal places.

    Project A:   %

    Project B:   %

  4. From your answers to parts a-c, which project would be selected?

    -Select-Project AProject BItem 7

    If the WACC was 18%, which project would be selected?

    -Select-Project AProject BItem 8

  5. Construct NPV profiles for Projects A and B. If an amount is zero, enter 0. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to the nearest cent.

    Discount Rate NPV Project A NPV Project B
    0% $        $       
    5
    10
    12
    15
    18.1
    23.33
  6. Calculate the crossover rate where the two projects' NPVs are equal. Do not round intermediate calculations. Round your answer to two decimal places.

      %

  7. What is each project's MIRR at a WACC of 18%? Do not round intermediate calculations. Round your answers to two decimal places.

    Project A:   %

    Project B:   %

appletImage?dbid=1990211259
Answer #1

Part a:

NPV of the two projects is calculated below:

Project A:

Year CF Discount Factor Discounted CF
0 $ -300.00 1/(1+0.12)^0= 1 1*-300= $ -300.00
1 $ -387.00 1/(1+0.12)^1= 0.892857143 0.892857142857143*-387= $ -345.54
2 $ -193.00 1/(1+0.12)^2= 0.797193878 0.79719387755102*-193= $ -153.86
3 $ -100.00 1/(1+0.12)^3= 0.711780248 0.711780247813411*-100= $    -71.18
4 $    600.00 1/(1+0.12)^4= 0.635518078 0.635518078404831*600= $    381.31
5 $    600.00 1/(1+0.12)^5= 0.567426856 0.567426855718599*600= $    340.46
6 $    850.00 1/(1+0.12)^6= 0.506631121 0.506631121177321*850= $    430.64
7 $ -180.00 1/(1+0.12)^7= 0.452349215 0.452349215336893*-180= $    -81.42
NPV = Sum of all Discounted CF $    200.41

Project B:

Year CF Discount Factor Discounted CF
0 $ -405.00 1/(1+0.12)^0= 1 1*-405= $ -405.00
1 $    132.00 1/(1+0.12)^1= 0.892857143 0.892857142857143*132= $    117.86
2 $    132.00 1/(1+0.12)^2= 0.797193878 0.79719387755102*132= $    105.23
3 $    132.00 1/(1+0.12)^3= 0.711780248 0.711780247813411*132= $      93.95
4 $    132.00 1/(1+0.12)^4= 0.635518078 0.635518078404831*132= $      83.89
5 $    132.00 1/(1+0.12)^5= 0.567426856 0.567426855718599*132= $      74.90
6 $    132.00 1/(1+0.12)^6= 0.506631121 0.506631121177321*132= $      66.88
NPV = Sum of all Discounted CF $    137.71

Part b: IRR is calculated either by hit and trial or by use of a financial calcualtor or Excel's goal seek function:

Project A: IRR = 18.10% rounded to 2 decimal places

Year CF Discount Factor 0 $ -300.00 1/(1+0.18096706596689)^0= 1 $ -387.00 1/(1+0.18096706596689)^1= 2 $ - 193.00 1/(1+0.18096

Project B: IRR comes to 23.33% rounded to 2 decimal places

Year CF Discount Factor 0 $ -405.00 1/(1+0.233310109087244)^0= 1 $ 132.00 1/(1+0.233310109087244)^1= 2 $ 132.00 1/(1+0.233310

Part c:

FVpositive CF -1 MIRR = 1 PVnegative CF

3679.27 MIRRA = V 952

MIRRA = 21.30%

MIRRB = 1199.75

MIRRR = 16.78%

Part D

NPV method reigns supreme above all so the project with highest NPV should be selected, which is project A. Even though the IRR of project B is higher.

NPV of the two projects if the WACC = 18% is given below:

Project A:

Year CF Discount Factor 0 $-300.00 1/(1+0.18)^0= 1 $ -387.00 1/(1+0.18)^1= 2 $ -193.00 1/(1+0.18)^2= 3 $ -100.00 1/(1+0.18)^3

Project B

Year CF Discount Factor 0 $ -405.00 1/(1+0.18)^0= 1 $ 132.00 1/(1+0.18)^1= 2 $ 132.00 1/(1+0.18)^2= 3 $ 132.00 1/(1+0.18)^3=

Even so Project A has a higher NPV and should be selected

NPV profile :

Rate A B
0% 2050 792
5% 1598.02 669.99
10% 1262.16 574.89
12% 1152.4 542.71
15% 1008.84 499.55
18% 882.85 460.5
23% 711.18 405.01
57.67% 213.99 213.99
80% 113.9 160.15
100% 69.53 129.94

2500 2000 1500 1000 500 213.99 0% 5% 10% 15% 20% 25% 30% 35% 40% 45% 50% 55% 60% 65% 70% 75% 80% 85% 90% 95% 100%105% AB

Part f) Crossover rate is 57.67%

Part g) MIRR when WACC is 18% is calculated below:

FVpositive CF -1 MIRR = 1 PVnegative CF

4830.54 MIRRA = V 883.95

MIRRA = 27.458%

MIRRR = 1470.68

MIRRR = 20.23%

So MIRR of both the projects increased